The value of A=2i+3j-4k and B= -2i+6j-7k then, what is the answer of (2/3AΒ  +Β  1/3B ) ?? Please give me answer of this question my mcqs test is on friday πŸ˜¨πŸ˜–

Grade:12

2 Answers

Arun
5 years ago
Dear student
Β 
2/3 A + 1/3 B = 1/3 (2A + B) = 1/3 ( 4i + 6j – 8k – 2i + 6j – 7k) = 1/3 (2i + 12 j – 15k)
Β 
Hope it helps
Β 
Regards
Arun
Khimraj
4 years ago
2*A + B =Β 2i + 12 j – 15k
1/3(2A+B) = = 2/3i + 4j – 5k
Hope it clears..........................................Β